OpenStudy (anonymous):

For the polynomial below, find F(-3). F(x)=-1/3x^2+8x-5

OpenStudy (cwrw238):

just plug in -3 for x and work it out

OpenStudy (anonymous):

Just change x to -3. \[f(-3)=-\frac{ (-3)² }{ 3 }+8*(-3)-5=-3-24-5=-32\]